OpenStudy (anonymous):

The table below shows the amounts of cherry juice and soda used in each container of cherry soda. The ratio of cherry juice to soda is 1:r, where r is the unit rate. Use the table to determine r. 1/4 4 2 2/3

OpenStudy (anonymous):



OpenStudy (anonymous):

would it be 4?

OpenStudy (ccswims):

yes, because 2/8 = 1/4 and 3/12 = 1/4

OpenStudy (ccswims):

so the ratio would be 1:4

OpenStudy (ccswims):

WAIT

OpenStudy (anonymous):

ok lol

OpenStudy (ccswims):

ok, it cant be 4 because 2 divided by 4 doesnt work

OpenStudy (anonymous):

the answer cant be a fraction right?

OpenStudy (ccswims):

ok, wait i confused myself there, it is 4. ok, im so sorry, idk why i made it so much more confusing than what it is. the ratio needs to be, for every 1 cup of cherry juice there are 4 cups of soda. so 4 would do that

OpenStudy (ccswims):

i kept thinking that i need to divide the fraction by the ratio, but the ratio itself is a simplified form of the fractions, so you dont do that

OpenStudy (ccswims):

so, yea it is B lol

OpenStudy (anonymous):

ohh ok so lets say if it was asking for 2 cups it would be 2:8 right?

OpenStudy (ccswims):

yea, every 2 cups of juice is 8 cups of soda. and every 4 cups of juice is 16 cups of soda

OpenStudy (ccswims):

as long as it stays to the 1:4 ratio its fine